# MicroLad: 2D-to-3D Microstructure Reconstruction and Generation via Latent Diffusion and Score Distillation

### v1

@@ 1 Introduction
Descriptor-based microstructure characterization and reconstruction (MCR) addresses this data gap by generating statistically representative microstructures that match prescribed statistical measures (e.g., n-point correlation functions), using methods such as the Yeong–Torquato algorithm [22, 23] and more recent gradient-based variants [24, 25, 26, 27, 28]. The core idea is to generate new, statistically representative microstructures from limited data, thereby augmenting datasets for analysis and simulation. This enables researchers to (i) produce large ensembles of equivalent microstructures, (ii) explore hypothetical “interpolated” states between known morphologies, and (iii) reconstruct full 3D volumes from sparse 2D observations. A major advantage of descriptor-based MCR is its physical interpretability: the same descriptors used as loss terms can be directly compared with experimental measurements. However, the approach is limited by the information content of the chosen descriptors. For instance, relying solely on two-point correlations may not accurately characterize complex microstructures [25], as distinct microstructural patterns may share identical or nearly identical two-point statistics. Consequently, multiple physically distinct structures can serve as equally valid "solutions" to the descriptor-matching problem
@@
Data-driven microstructure reconstruction approaches, by contrast, learn microstructure distributions directly from images. Widely used architectures include deep generative models such as variational autoencoders (VAEs) [29, 30, 31, 32, 33, 34] and generative adversarial networks (GANs) [35, 36, 37, 38, 39, 40]. For example, SliceGAN [36], a GAN-based framework, has demonstrated 2-D-to-3-D reconstruction for a variety of anisotropic and multiphase microstructures. Instead of manually defining descriptors, these models learn salient morphological features from the data and can synthesise new instances that remain statistically close to the training set. More recently, diffusion models have emerged as state-of-the-art generative frameworks, achieving remarkable performance across diverse tasks due to their high fidelity, diversity, and stable training dynamics [41, 42, 43, 44, 45, 46, 47]. These models operate by progressively corrupting and then recovering data through a predefined forward and reverse diffusion process, typically formulated as a Markovian process—as in denoising diffusion probabilistic models (DDPMs) [42]. In addition, latent diffusion models (LDMs) [48] perform the diffusion steps in a learned, lower-dimensional latent space, reducing the computational cost while preserving essential semantic and structural features. This allows them to generate samples that faithfully follow the underlying data manifold. Several studies have shown that 2D diffusion models can faithfully reproduce complex microstructures—spherical inclusions, ceramics, composites, grain structures, and more [49, 50]. Furthermore, to address the 2D-to-3D reconstruction challenge, Lee and Yun proposed a multi-plane denoising diffusion (MPDD) framework [51, 52]. Their approach denoises multiple orthogonal 2D planes simultaneously with a pretrained 2D diffusion model, enforcing spatial coherence across slices and thereby synthesising statistically equivalent 3D microstructures.
@@
Figure 1: MicroLad framework description: (a) acquire a 2D microstructure dataset, (b) encode the data into a latent space, (c) add noise to the latent variables, (d) train a U-Net to predict the noise, (e) initialize stacked 2D latent variables, (f) apply L-MPDD to spatially connect the 2D reverse diffusion processes, (g) generate 3D latent variables, (h) refine the latent representation to produce 3D voxels, (i) decode the final 3D microstructure, (j) apply SDS for inverse-controlled microstructure generation with different objectives, and (k) generate 3D microstructures by targeting different design objectives.
@@ 2.2 Dimensionality expansion (2D-to-3D) for microstructure data
$\mathbf{x}_{2D}\in\mathcal{X}_{2D},\quad\mathbf{x}_{2D}\subseteq\mathbb{R}^{H\times W},$ (1)
where $\mathbf{x}_{2D}$ denotes the experimental 2D microstructure data, and $H$ and $W$ represent its spatial dimensions (height and width, respectively). To model this data distribution, we train a 2D generative model $p_{\theta}$, that approximates the true but unknown distribution $p_{\text{data}}$ of experimental microstructures:
$p_{\theta}(\mathbf{x}_{2D})\approx p_{\text{data}}(\mathbf{x}_{2D}),$ (2)
@@
$\hat{\mathbf{x}}_{2D}^{(i)}\sim p_{\theta}(\mathbf{x}_{2D}),\quad i=1,2,\dots,D,$ (3)
where $\hat{\mathbf{x}}_{2D}^{(i)}$ is the $i$-th generated 2D microstructure slice and $D$ denotes the total number of slices required to form a 3D microstructure. In other words, we can think of 2D-to-3D microstructure reconstruction as generating multiple 2D slices that are spatially connected and collectively represent a 3D microstructure volume. In this context, the goal is now to generate multiple 2D slices $\{\hat{\mathbf{x}}_{2D}^{(i)}\}_{i=1}^{D}$ that are spatially coherent and collectively represent a volume $\mathbf{x}_{3D}$:
$\mathbf{x}_{3D}\in\mathcal{X}_{3D},\quad\mathbf{x}_{3D}\subseteq\mathbb{R}^{H\times W\times D},$ (4)
@@
$\hat{\mathbf{x}}_{2D}^{(i)}\sim p_{\theta}(\mathbf{x}_{2D}\mid\hat{\mathbf{x}}_{2D}^{(j)}),\quad j\in\mathcal{N}(i),$ (5)
@@ 2.3 Inverse-controlled 2D-to-3D microstructure generation
$\mathbf{x}_{3D}^{*}=\underset{\mathbf{x}_{3D}\in\mathcal{X}_{3D}}{\arg\min}\;\left[\mathcal{L}_{M}(\mathbf{M}(\mathbf{x}_{3D}),\mathbf{M}^{*})+\mathcal{L}_{P}(\mathbf{P}(\mathbf{x}_{3D}),\mathbf{P}^{*})\right],$ (11)
subject to the critical constraint (5) that the 3D microstructure ($\mathbf{x}_{3D}$) must be generated from spatially connected 2D slices ($\hat{\mathbf{x}}_{2D}^{(i)}$) sampled from the trained $p_{\theta}$.
@@ 3.1 Denoising diffusion probabilistic models (DDPMs)
$\mathcal{L}_{\text{DDPM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\epsilon}[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}],$ (15)
where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is Gaussian noise, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$. The sampling (generation) procedure involves iteratively applying the learned reverse denoising step from pure Gaussian noise:
@@ 3.2 Latent diffusion models (LDMs)
where $C$, $H^{\prime}$, and $W^{\prime}$ represent latent space dimensions. The forward diffusion process in latent space is defined as:
@@
The reverse generative process in latent space, parameterized by $\boldsymbol{\varepsilon}_{\theta}$, is similarly defined:
@@
$\mathcal{L}_{\text{LDM}}(\theta)=\mathbb{E}_{t,\mathbf{z}_{0},\boldsymbol{\varepsilon}}[\|\boldsymbol{\varepsilon}-\boldsymbol{\varepsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\varepsilon},t)\|^{2}],$ (21)
where $\boldsymbol{\varepsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Finally, to generate new samples, latent variables are sampled through reverse diffusion and decoded back into data space using a decoder $\mathcal{D}$:
@@
According to Lee and Yun [52, 51], MPDD—which performs iterative 2D denoising diffusion along the three Cartesian axes of a 3D volume—ensures that 2D denoising steps maintain spatial coherence in 3D. In this study, we adopt a similar approach but implement MPDD within the latent space (i.e., L-MPDD) using a pretrained 2D LDM. Because the pretrained 2D LDM lacks an explicit mechanism for enforcing 3D connectivity (5), we leverage L-MPDD to implicitly achieve this goal by applying denoising diffusion along all three orthogonal latent dimensions—height $H^{\prime}$, width $W^{\prime}$, and depth $D^{\prime}$.
@@
Each denoising step applies the pretrained 2D diffusion model $p_{\theta}$, independently across slices, thereby implicitly enforcing 3D spatial consistency. After completing the denoising process, we obtain the latent representation:
@@
Due to the mismatch between the latent volume shape $\mathbf{Z}\in\mathbb{R}^{D^{\prime}\times C\times H^{\prime}\times W^{\prime}}$ and the desired final microstructure resolution $\mathbf{x}_{3D}\in\mathbb{R}^{D\times H\times W}$, we apply a refinement step to improve spatial coherence and fidelity. We initialize an empty volume at the final resolution:
@@
After $K$ refinement iterations, the final 3D microstructure is obtained:
$\mathbf{x}_{3D}=\mathbf{X}^{(K)}\in\mathbb{R}^{D\times H\times W}.$ (36)
@@ 3.4 Inverse-controlled 2D-to-3D microstructure generation with score-distillation sampling
After reconstructing the 3D microstructure volume $\mathbf{x}_{3D}\in\mathbb{R}^{D\times H\times W}$, we apply SDS to refine randomly selected 2D slices within the volume, guiding the 3D microstructure toward user-specified target microstructural descriptors ($\mathbf{M}^{*}$) and target effective material properties ($\mathbf{P}^{*}$). At each SDS iteration, a 2D slice $\mathbf{x}_{\text{slice}}$ is randomly selected from the 3D volume along one of the three Cartesian axes (depth, height, or width). The selected slice is then encoded into the latent space:
@@
$\mathcal{L}_{\text{SDS}}=\left\|\varepsilon-\varepsilon_{\theta}\left(\mathbf{z}_{\text{slice},t},t\right)\right\|^{2}.$ (39)
@@
To assess material property accuracy, a differentiable physics solver (e.g., differentiable FEM $\mathcal{H}$) is used:
@@
The total loss is:
$\mathcal{L}_{\text{total}}$ $=\mathcal{L}_{\text{SDS}}+w_{M}\mathcal{L}_{M}+w_{P}\mathcal{L}_{P},$ (43)
$\mathbf{z}_{\text{slice}}$ $\leftarrow\mathbf{z}_{\text{slice}}-\eta\nabla_{\mathbf{z}_{\text{slice}}}\mathcal{L}_{\text{total}},$ (44)
where $w_{M}$ and $w_{P}$ are weighting terms, and $\eta$ is the learning rate. The updated latent slice is decoded and replaces the original:
@@ 3.5 Implementation details
The VAE was pretrained on a variety of microstructure types—including Voronoi grains, generic grain microstructures, spherical inclusions, aligned elliptical inclusions, graphite, nickel-based superalloys (NBSA), Gaussian random fields (GRF), and random elliptical inclusions—using the open-source microstructure dataset from Robertson et al.[59]. It was then fine-tuned on the target microstructures of interest: binary carbonate and three-phase SOFC microstructures. Latent representations were encoded as $4\times 16\times 16$ feature maps. The LDM operates on this latent space and was trained for 1000 diffusion steps using a linear noise schedule $\beta\in[1\times 10^{-4},2\times 10^{-2}]$. Full architectural specifications and hyperparameters for both the VAE and LDM are provided in Appendix A.
@@
SDS is then applied iteratively to optimize randomly selected 2D slices from the reconstructed 3D volume (Section 3.4). Each latent slice is updated using a gradient-based loss that combines (i) a diffusion score loss to maintain realism, (ii) descriptor-based losses for matching volume fraction or surface area, and (iii) a differentiable physical loss based on effective diffusivity computed via FEM. Optimization is performed in the latent space using the Adam optimizer, with predefined SDS learning rate and number of SDS steps.

### v3

@@ 1 Introduction
Descriptor-based microstructure characterization and reconstruction (MCR) addresses this data gap by generating statistically representative microstructures that match prescribed statistical measures (e.g., n-point correlation functions), using methods such as the Yeong–Torquato algorithm [22, 23] and more recent gradient-based variants [24, 25, 26, 27, 28]. The core idea is to generate new, statistically representative microstructures from limited data, thereby augmenting datasets for analysis and simulation. This enables researchers to (i) produce large ensembles of equivalent microstructures, (ii) explore hypothetical “interpolated” states between known morphologies, and (iii) reconstruct full 3D volumes from sparse 2D observations. A major advantage of descriptor-based MCR is its physical interpretability: the same descriptors used as loss terms can be directly compared with experimental measurements. However, the approach is limited by the information content of the chosen descriptors. For instance, relying solely on two-point correlations may not accurately characterize complex microstructures [25], as distinct microstructural patterns may share identical or nearly identical two-point statistics. Consequently, multiple physically distinct structures can serve as equally valid "solutions" to the descriptor-matching problem.
@@
Data-driven microstructure reconstruction approaches, by contrast, learn microstructure distributions directly from images. Widely used architectures include deep generative models such as variational autoencoders (VAEs) [29, 30, 31, 32, 33, 34] and generative adversarial networks (GANs) [35, 36, 37, 38, 39, 40]. For example, SliceGAN [36], a GAN-based framework, has demonstrated 2D-to-3D reconstruction for a variety of anisotropic and multiphase microstructures. Instead of manually defining descriptors, these models learn salient morphological features from the data and can synthesize new instances that remain statistically close to the training set. More recently, diffusion models have emerged as state-of-the-art generative frameworks, achieving remarkable performance across diverse tasks due to their high fidelity, diversity, and stable training dynamics [41, 42, 43, 44, 45, 46, 47]. These models operate by progressively corrupting and then recovering data through a predefined forward and reverse diffusion process, typically formulated as a Markovian process—as in denoising diffusion probabilistic models (DDPMs) [42]. In addition, latent diffusion models (LDMs) [48] perform the diffusion steps in a learned, lower-dimensional latent space, reducing the computational cost while preserving essential semantic and structural features. This allows them to generate samples that faithfully follow the underlying data manifold. Several studies have shown that 2D diffusion models can faithfully reproduce complex microstructures—spherical inclusions, ceramics, composites, grain structures, and more [49, 50]. Furthermore, to address the 2D-to-3D reconstruction challenge, Lee and Yun proposed a multi-plane denoising diffusion (MPDD) framework [51, 52]. Their approach denoises multiple orthogonal 2D planes simultaneously with a pretrained 2D diffusion model, enforcing spatial coherence across slices and thereby synthesizing statistically equivalent 3D microstructures.
@@
Figure 1: MicroLad framework description: (a) acquire a 2D microstructure dataset, (b) encode the data into a latent space, (c) add noise to the latent variables, (d) train a U-Net to predict the noise, (e) initialize stacked 2D latent variables, (f) apply L-MPDD to spatially connect the 2D reverse diffusion processes, (g) generate 3D latent variables, (h) refine the latent representation to produce 3D voxels, (i) decode the final 3D microstructure, and (j) apply SDS for inverse-controlled microstructure generation with different objectives.
@@ 2.2 Dimensionality expansion (2D-to-3D) for microstructure data
$\mathbf{x}_{\mathrm{2D}}\in\mathcal{X}_{\mathrm{2D}},\quad\mathbf{x}_{\mathrm{2D}}\subseteq\mathbb{R}^{H\times W},$ (1)
where $\mathbf{x}_{\mathrm{2D}}$ denotes the experimental 2D microstructure data, and $H$ and $W$ represent its spatial dimensions (height and width, respectively). To model this data distribution, we train a 2D generative model $p_{\theta}$, that approximates the true but unknown distribution $p_{\text{data}}$ of experimental microstructures:
$p_{\theta}(\mathbf{x}_{\mathrm{2D}})\approx p_{\text{data}}(\mathbf{x}_{\mathrm{2D}}),$ (2)
@@
$\hat{\mathbf{x}}_{\mathrm{2D}}^{(i)}\sim p_{\theta}(\mathbf{x}_{\mathrm{2D}}),\quad i=1,2,\dots,D,$ (3)
where $\hat{\mathbf{x}}_{\mathrm{2D}}^{(i)}$ denotes the $i$-th generated 2D microstructure slice, and $D$ represents the total number of slices required to construct a 3D microstructure along the depth direction. In other words, we can think of 2D-to-3D microstructure reconstruction as generating multiple 2D slices that are spatially connected and collectively represent a 3D microstructure volume. In this context, the goal is now to generate multiple 2D slices $\{\hat{\mathbf{x}}_{\mathrm{2D}}^{(i)}\}_{i=1}^{D}$ that are spatially coherent and collectively represent a volume $\mathbf{x}_{\mathrm{3D}}$:
$\mathbf{x}_{\mathrm{3D}}\in\mathcal{X}_{\mathrm{3D}},\quad\mathbf{x}_{\mathrm{3D}}\subseteq\mathbb{R}^{H\times W\times D},$ (4)
@@
$\hat{\mathbf{x}}_{\mathrm{2D}}^{(i)}\sim p_{\theta}(\mathbf{x}_{\mathrm{2D}}\mid\hat{\mathbf{x}}_{\mathrm{2D}}^{(j)}),\quad j\in\mathcal{N}(i),$ (5)
@@ 2.3 Inverse-controlled 2D-to-3D microstructure generation
$\mathbf{x}_{\mathrm{3D}}^{*}=\underset{\mathbf{x}_{\mathrm{3D}}\in\mathcal{X}_{\mathrm{3D}}}{\arg\min}\;\left[\mathcal{L}_{M}(\mathbf{M}(\mathbf{x}_{\mathrm{3D}}),\mathbf{M}^{*})+\mathcal{L}_{P}(\mathbf{P}(\mathbf{x}_{\mathrm{3D}}),\mathbf{P}^{*})\right],$ (11)
subject to the critical constraint (5) that the 3D microstructure ($\mathbf{x}_{\mathrm{3D}}$) must be generated from spatially connected 2D slices ($\hat{\mathbf{x}}_{\mathrm{2D}}^{(i)}$) sampled from the trained $p_{\theta}$.
@@ 3.1 Denoising diffusion probabilistic models (DDPMs)
$\mathcal{L}_{\text{DDPM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\varepsilon}[\|\varepsilon-\varepsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\varepsilon,t)\|^{2}],$ (15)
where $\varepsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is Gaussian noise, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$. The sampling (generation) procedure involves iteratively applying the learned reverse denoising step from pure Gaussian noise:
@@ 3.2 Latent diffusion models (LDMs)
here $C$, $H^{\prime}$, and $W^{\prime}$ represent latent space dimensions. The forward diffusion process in latent space is defined as:
@@
The reverse generative process in latent space, parameterized by ${\varepsilon}_{\theta}$, is similarly defined:
@@
$\mathcal{L}_{\text{LDM}}(\theta)=\mathbb{E}_{t,\mathbf{z}_{0},{\varepsilon}}[\|{\varepsilon}-{\varepsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\varepsilon},t)\|^{2}],$ (21)
where ${\varepsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Finally, to generate new samples, latent variables are sampled through reverse diffusion and decoded back into data space using a decoder $\mathcal{D}$:
@@
According to Lee and Yun [52, 51], MPDD (which performs iterative 2D denoising diffusion along the three Cartesian axes of a 3D volume) effectively approximates the 3D reverse diffusion process. In this study, we adopt a similar approach but implement MPDD within the latent space (i.e., L-MPDD) using a pretrained 2D LDM. Because the pretrained 2D LDM lacks an explicit mechanism for enforcing 3D connectivity (5), we leverage L-MPDD to implicitly achieve this goal by applying denoising diffusion along all three orthogonal latent dimensions: height $H^{\prime}$, width $W^{\prime}$, and depth $D^{\prime}$.
@@
Each denoising step applies the trained 2D diffusion model $p_{\theta}$ across multiple planes, thereby implicitly enforcing 3D spatial consistency. After completing the denoising process, we obtain the latent representation:
@@
Due to the mismatch between the latent volume shape $\mathbf{Z}\in\mathbb{R}^{C\times D^{\prime}\times H^{\prime}\times W^{\prime}}$ and the desired final microstructure resolution $\mathbf{x}_{\mathrm{3D}}\in\mathbb{R}^{D\times H\times W}$, we apply a refinement step to improve spatial coherence and fidelity. We initialize an empty volume at the final resolution:
@@
After $K$ refinement iterations, the final 3D microstructure is obtained (additional schematic in Appendix A):
$\mathbf{x}_{\mathrm{3D}}=\mathbf{X}^{(K)}\in\mathbb{R}^{D\times H\times W}.$ (36)
@@ 3.4 Inverse-controlled 2D-to-3D microstructure generation with score-distillation sampling
After reconstructing the 3D microstructure volume $\mathbf{x}_{\mathrm{3D}}\in\mathbb{R}^{D\times H\times W}$, we apply SDS to refine randomly selected 2D slices within the volume, guiding the 3D microstructure toward user-specified target microstructural descriptors ($\mathbf{M}^{*}$) and target effective material properties ($\mathbf{P}^{*}$). At each SDS iteration, a 2D slice $\mathbf{x}_{\text{slice}}$ is randomly selected from the 3D volume along one of the three Cartesian axes (depth, height, or width). The selected slice is then encoded into the latent space:
@@
$\mathcal{L}_{\text{SDS}}=\kappa(t)\,\|\varepsilon-\varepsilon_{\theta}(\mathbf{z}_{\text{slice},t},t)\|^{2},$ (39)
$\kappa(t)=\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}.$ (40)
@@
To define the objective function for property-guided generation, a differentiable physics solver (e.g., differentiable FEM, $\mathcal{H}$) is employed, and the corresponding loss is formulated as follows:
@@
Then, we update the latent slice as follows:
$\mathbf{z}_{\text{slice}}\;\leftarrow\;\mathbf{z}_{\text{slice}}-\eta\Big(\nabla^{\mathrm{pg}}_{\mathbf{z}_{\text{slice}}}\mathcal{L}_{\text{SDS}}+w_{M}\,\nabla_{\mathbf{z}_{\text{slice}}}\mathcal{L}_{M}+w_{P}\,\nabla_{\mathbf{z}_{\text{slice}}}\mathcal{L}_{P}\Big),$ (44)
where $w_{M}$ and $w_{P}$ are weighting terms and $\eta$ is the learning rate. Here $\nabla^{\mathrm{pg}}$ denotes a pseudo-gradient update. In the SDS term, we treat $\varepsilon_{\theta}(\cdot)$ as a constant target and do not back-propagate through the frozen diffusion network (i.e., we drop the Jacobian $\partial\varepsilon_{\theta}/\partial\mathbf{z}_{\text{slice}}$). This mirrors the original SDS setup [53], where the diffusion prior is frozen and only the external generator is optimized. In our case, the “external generator” is the latent slice itself. The updated latent slice is decoded and replaces the original:
@@ 3.5 Implementation details
The VAE was pretrained on a variety of microstructure types—including Voronoi grains, generic grain microstructures, spherical inclusions, aligned elliptical inclusions, graphite, nickel-based superalloys (NBSA), Gaussian random fields (GRF), and random elliptical inclusions—using the open-source microstructure dataset from Robertson et al.[59]. It was then fine-tuned on the target microstructures of interest: binary carbonate and three-phase SOFC microstructures. Latent representations were encoded as $4\times 16\times 16$ feature maps. The LDM operates on this latent space and was trained for 1000 diffusion steps using a linear noise schedule $\beta\in[1\times 10^{-4},2\times 10^{-2}]$. Full architectural specifications and hyperparameters for both the VAE and LDM are provided in Appendix B.
@@
SDS is then applied iteratively to optimize randomly selected 2D slices from the reconstructed 3D volume (Section 3.4). Each latent slice is updated using a gradient-based loss that combines (i) a diffusion score loss to maintain realism, (ii) descriptor-based losses for matching volume fraction or surface area, and (iii) a differentiable physical loss based on effective diffusivity computed via FEM.